# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Appendix D Details of the computer algebra for general spin-1/2 systems
We discuss the details of the algorithm based on computer algebra and analyze the computational time. We focus entirely on spin-1/2 systems for concreteness, while a similar analysis should be possible for spin systems with any spin quantum number and fermionic systems.
@@
We propose a computer algebra framework of spin-1/2 systems. This framework inherits basic concepts from previous general frameworks of computer algebra [72, 73, 74, 68, 33] but is more efficient and suited for our purpose.
@@
is an operator composed of three terms. We use $\sigma,\tau,\dots$ to denote tensor products of Pauli operators, including $I$ as a special case, and we use $z_{A,\sigma}$ to denote the scalar coefficient before $\sigma$ in $A$. For example, $z_{A,\sigma}=1.5$ for $\sigma=X_{1}Z_{3}$ in Eq. (62). We define $|A|$ as the number of terms in $A$, e.g., $|A|=3$ in the example in Eq. (62). Below, we assume that all terms are $k$-local, with $k$ small and independent of the system size. Based on this assumption, we will disregard the number of Pauli operators per term and only consider the number of terms to evaluate the computational time.
Our computer algebra framework stores an operator using a data structure called a hash table [81]. A hash table is a set of correspondences from a “key” to a “value,” which, in our case, is from a tensor product $\sigma$ to the scalar coefficient $z_{A,\sigma}$. A tensor product of Pauli operators is represented by a list of Pauli operator kinds and indices, sorted in the ascending order of the spin index. For example, $X_{1}Z_{3}$ is represented by a list $[(X,1),(Z,3)]$, and the entire operator $A$ in Eq. (62) is stored as
[…]
@@
A hash table allows a quick lookup of its elements [81]. Given an arbitrary $\sigma$ and a hash-table representation of an arbitrary operator $A$, it takes only $O(1)$ time to determine whether $\sigma$ is contained in $A$ and, if so, to obtain the scalar coefficient $z_{A,\sigma}$. This lookup time is independent of $|A|$ for ideal hash tables.
@@
The multiplication of two operators, $A$ and $B$, is completed in $O(|A||B|)$ time by the following procedure. We prepare an empty hash table for storing the result. We pick up one term $\sigma\mapsto z_{A,\sigma}$ from $A$ and one term $\tau\mapsto z_{B,\tau}$ from $B$, multiply the two tensor products $\sigma$ and $\tau$ according to the rules such as $X_{i}Y_{i}=iZ_{i}$, $Y_{i}Z_{i}=iX_{i},$ $Z_{i}X_{i}=iY_{i}$ and $X_{i}X_{i}=Y_{i}Y_{i}=Z_{i}Z_{i}=I_{i}$, multiply their scalar coefficients, and insert $\sigma\tau\mapsto z_{A,\sigma}z_{B,\tau}$ into the hash table for storing the result. This procedure for each pair of terms is done in $O(1)$ time. We repeat this procedure for all pairs of terms. Since there are $|A||B|$ pairs, the total computation time is $O(|A||B|)$. This computation time is also expressed as $O(|AB|)$ because the resulting product $AB$ usually contains $|AB|=O(|A||B|)$ terms.
@@
We can efficiently compute traces and commutators by developing special algorithms that take less time than naive approaches. These algorithms will be crucial for efficiently calculating weighted actions.
The trace $\mathop{\mathrm{tr}}(AB)$ for two operators $A$ and $B$ can be computed in $O(\min\{|A|,|B|\})$ time, which is faster than computing the product $AB$. To achieve this performance, we use the fact that $\mathop{\mathrm{tr}}(\sigma)=0$ for any tensor product of Pauli operators $\sigma\neq I$, except that $\mathop{\mathrm{tr}}(I)=2^{N}$. Moreover, the product $\sigma\tau$ becomes the identity operator if and only if $\sigma=\tau$. Therefore, the trace $\mathop{\mathrm{tr}}(AB)$ is given by $\mathop{\mathrm{tr}}(AB)=2^{N}\sum_{\sigma}z_{A,\sigma}z_{B,\sigma}$, where the sum is over all $\sigma$’s that appear in both $A$ and $B$. If $|A|\leq|B|$, we can calculate $2^{-N}\mathop{\mathrm{tr}}(AB)$ by picking up each term $\sigma\mapsto z_{A,\sigma}$ in $A$, looking up whether $\sigma$ appears in $B$, and if so, adding $z_{A,\sigma}z_{B,\sigma}$ to the result. The single lookup process is done in $O(1)$ time regardless of $|B|$ using a hash table, and thus, the total computational time is $O(|A|)$. If $|A|\geq|B|$, we can swap the roles of $A$ and $B$ to finish the calculation in $O(|B|)$ time. Therefore, we can compute $\mathop{\mathrm{tr}}(AB)$ in $O(\min\{|A|,|B|\})$ time.
@@
The commutator $[B,A]$ for two operators $B$ and $A$ also has a fast algorithm. The commutator of two terms is zero if they have no spins in common, e.g., $[X_{1}Y_{2}Z_{3},X_{4}Y_{5}]=0$. Therefore, we do not need to iterate through all pairs of terms of $B$ and $A$, but we can skip most of them. This shortcut is achieved by the following procedure. First, we assign integers $1,2,\dots,|A|$ to the terms of $A$ in an arbitrary order, and we prepare the set of indices $\Theta_{i}(A)=\{l\mid\text{$l$th term of $A$ acts on spin $i$}\}$ for $i=1,\dots,N$. This is done by iterating through the terms of $A$ once, taking $O(|A|)$ time. Next, we prepare an empty hash table to store the result of $[B,A]$. We iterate through each term $\sigma\mapsto z_{B,\sigma}$ of $B$. If, for example, $\sigma$ acts on spins $i=3,4$, we compute the commutator between $\sigma$ and the terms of $A$ whose indices are in $\Theta_{3}(A)\cup\Theta_{4}(A)$, and we add the result to the hash table representing the result.
This procedure of computing $[B,A]$ runs in $O(N^{-1}|A||B|)$ time under the following assumptions in addition to the $k$-local assumption for all operators. We assume that $A$ consists of $O(N)$ or more terms, and the terms are spread evenly over an $O(N)$ number of spins so that each spin is acted by at most $O(N^{-1}|A|)$ terms. Thus, the size of the index set $|\Theta_{i}(A)|$ is at most $O(N^{-1}|A|)$. For every term in $B$, we compute the commutator with the union of at most $k$ sets out of $\Theta_{1}(A),\dots,\Theta_{N}(A)$ because $B$ is a $k$-local operator. The union contains at most $O(N^{-1}|A|)$ terms, where we neglect $k$ because $k$ is assumed to be small. Thus, the iteration through $B$ takes $O(N^{-1}|A||B|)$ time in total. This analysis also shows that the resulting commutator has $|[B,A]|=O(N^{-1}|A||B|)$ terms.
By extending this algorithm, we can also efficiently compute the commutators $[B,A_{\mu}]$ between an operator $B$ and a series of operators $A_{1},A_{2},\dots$. To do so, we prepare the set $\Theta_{i}(\{A_{\mu}\})=\{(\mu,l)\mid\text{$l$th term of $A_{\mu}$ acts on spin $i$}\}$, and we compute the commutators $[B,A_{1}],[B,A_{2}],\ldots$ simultaneously by iterating through the terms of $B$ once, similarly to the above procedure. The computation time is estimated by replacing $|A|$ with $\sum_{\mu}|A_{\mu}|$ in the previous paragraph. Assuming that $A_{1},A_{2},\dots$ have $O(N)$ or more terms in total, and the terms spread over $O(N)$ number of spins without concentrating on a few spins, the total computation takes $O(N^{-1}|B|\sum_{\mu}|A_{\mu}|)$ time, and the resulting number of terms is $\sum_{\mu}|[B,A_{\mu}]|=O(N^{-1}|B|\sum_{\mu}|A_{\mu}|)$ in total.
@@
We detail the entire algorithm for computing the optimal driving coefficients $\bm{\alpha}^{(\mathcal{P})}(\lambda)$ in the general framework in Sec. II, and we show that the computation time scales as $O(N^{K})$ with respect to the system size $N$ and degree $K$. The algorithms are designed to take full advantage of computer algebra. We use notation from Appendices D.1 and D.2: we use $|A|$ for the number of terms in $A$, and $\mathop{\mathrm{tr}}(A*BC)$ to emphasize that we evaluate the trace by preparing $A$ and $BC$ without computing the entire product $ABC$.
@@
[$\partial{}_{\lambda}[H(\lambda)^{k}]=\sum_{l=0}^{k-1}H(\lambda)^{l}[\partial_{\lambda}H(\lambda)]H(\lambda)^{k-1-l}.$ (64)
@@
(3) Compute the polynomial $\mathcal{P}_{\lambda}(H(\lambda))$ and its derivative $\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))$,
$\mathcal{P}_{\lambda}(H(\lambda))$ $=\sum_{k=1}^{K}p_{k}(\lambda)H(\lambda)^{k},$ (65)
$\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))$ $=\sum_{k=1}^{K}p_{k}(\lambda)\partial{}_{\lambda}[H(\lambda)^{k}],$ (66)
@@
(4) Compute the commutators $[\mathcal{P}(H(\lambda)),A_{\mu}]$ for all $\mu$. Since $|\mathcal{P}(H(\lambda))|=O(N^{K})$ and $\sum_{\mu}|A_{\mu}|=O(N)$, the computation time is $O(N^{-1}NN^{K})=O(N^{K})$, and the resulting commutators contain $O(N^{K})$ terms in total.
(5) Compute the double commutators $[[\mathcal{P}(H(\lambda)),A_{\mu}],A_{\nu}]$ for all $\mu\leq\nu$. Using $\sum_{\mu}|[\mathcal{P}(H(\lambda)),A_{\mu}]|=O(N^{K})$ and $\sum_{\mu}|A_{\mu}|=O(N)$, the total computation time is $\sum_{\mu}O(N^{-1}|[\mathcal{P}(H(\lambda)),A_{\mu}]|\sum_{\nu}|A_{\nu}|)=O(N^{-1}N^{K}N)=O(N^{K})$, and the total number of the resulting terms is $O(N^{K})$.
(6) Compute $r_{\mu}^{(K)}(\lambda)$ for all $\mu$ by
$r_{\mu}^{(\mathcal{P})}(\lambda)=i\hbar\,\mathop{\mathrm{tr}}\left\{\partial^{\prime}_{\lambda}\mathcal{P}(H(\lambda))*[\mathcal{P}(H(\lambda)),A_{\mu}]\right\}.$ (67)
Since $|[\mathcal{P}(H(\lambda)),A_{\mu}]|$ is less than $|\partial^{\prime}_{\lambda}\mathcal{P}(H(\lambda))|$, the calculation of $r_{\mu}^{(\mathcal{P})}(\lambda)$ for a single $\mu$ is done in $O\bigl{(}|[\mathcal{P}(H(\lambda)),A_{\mu}]|\bigr{)}$ time. Thus, the calculation for all $\mu$ takes the time $O\bigl{(}\sum_{\mu}|[\mathcal{P}(H(\lambda)),A_{\mu}]|\bigr{)}=O(N^{K})$.
(7) Compute $Q_{\mu\nu}^{(K)}(\lambda)$ for all $\mu\leq\nu$ by
$Q_{\mu\nu}^{(\mathcal{P})}(\lambda)=\mathop{\mathrm{tr}}\left\{\mathcal{P}(H(\lambda))*[[\mathcal{P}(H(\lambda)),A_{\mu}],A_{\nu}]\right\},$ (68)
where we use $-\mathop{\mathrm{tr}}([A,B]C)=\mathop{\mathrm{tr}}(A[C,B])$ to deduce Eq. (68) from Eq. (17a). The calculation for all $\mu\leq\nu$ takes the time $O\bigl{(}\sum_{\mu\leq\nu}|[[\mathcal{P}(H(\lambda)),A_{\mu}],A_{\nu}]|\bigr{)}=O(N^{K})$. Note that $Q_{\mu\nu}^{(K)}(\lambda)$ for $\mu>\nu$ is known from the symmetry $Q_{\mu\nu}^{(K)}(\lambda)=Q_{\nu\mu}^{(K)}(\lambda)$.
(8) Solve the $M$-variate linear equation in Eq. (18) to determine $\bm{\alpha}^{(\mathcal{P})}(\lambda)$ and output $\bm{\alpha}^{(\mathcal{P})}(\lambda)$. Although the standard numerical method takes $O(N^{3})$ time, we do not take this scaling into account because this step requires no algebraic manipulations and is usually much faster than the preceding steps.
In total, the computation time scales as $O(N^{K})$ if we neglect the time for step (8).
Algorithm 2.—This algorithm assumes that the Hamiltonian has the form of Eq. (19). We use symbols and definitions in Appendix B.3.
@@
(1) Compute the operators $\tilde{F}_{g}$ for $g=1,2,\dots$ by multiplication and addition of $F_{\gamma}$’s.
(2) For $g=1,2,\dots$, compute the commutators $[\tilde{F}_{g},A_{\mu}]$ for $\mu=1,\dots,M$ and the double commutators $[[\tilde{F}_{g},A_{\mu}],A_{\nu}]$ for all $\mu\leq\nu$.
(3) For all $g,g^{\prime}=1,2,\dots$, compute
@@
(4) Compute the scalar coefficient $\tilde{f}_{g}(\lambda)$ for $g=1,2,\dots$ by multiplying $f_{\gamma}(\lambda)$’s, and compute $\partial_{\lambda}\tilde{f}_{g}(\lambda)$ for $g=1,2,\dots$ using the Leibniz rule.
(5) Compute $r_{\mu}^{(\mathcal{P})}(\lambda)$ for all $\mu$ and $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ for all $\mu\leq\nu$ by Eq. (58). Note that $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ for $\mu>\nu$ is known from the symmetry $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)=Q_{\nu\mu}^{(\mathcal{P})}(\lambda)$.
@@
The computation time of Algorithm 2 is dominated by the first stage. Although the second stage is repeated many times, it is usually much faster than the first stage because it involves only scalar computations and does not require any algebraic computations. The computation time for the steps in the first stage is the same as for the corresponding steps in Algorithm 1. Concretely, step (1) in Algorithm 2 corresponds to step (1) in Algorithm 1, (2) corresponds to (4) and (5), and (3) corresponds to (6) and (7). Thus, the total computation time scales as $O(N^{K})$.
@@
This subsection presents entire algorithms specialized for ground-state evolution in Sec. III. These algorithms combine the algorithms in Appendix D.3 with the minimization of $\Omega_{\lambda}^{(K)}(E)$ discussed in Appendices C.1 and C.2. They are slightly simplified by taking advantage of the simple form of the polynomial $\mathcal{P}^{\mathrm{GS},K}(H(\lambda))=[H(\lambda)-E_{\lambda}^{(K)}]^{K}$. As in Appendix D.3, we present two algorithms, algorithm 3 for a single value of $\lambda$ and algorithm 4 for an entire range of $\lambda$. Our C++ code on GitHub [69] implements algorithm 3 and algorithm 4.
Algorithm 3.—This algorithm takes the algebraic representations of $H(\lambda)$, $\partial_{\lambda}H(\lambda)$, and $A_{\mu}$ for $\mu=1,\dots,M$ as input. The algorithm proceeds as follows:
@@
(3) Minimize the function $\Omega_{\lambda}^{(K)}(E)$ to determine $E_{\lambda}^{(K)}$. The details of the minimization are discussed in Appendix C.1.
(4) Compute $[H(\lambda)-E_{\lambda}^{(K)}]^{k}$ for $k=2,\dots,K$ by
$[H(\lambda)-E_{\lambda}^{(K)}]^{k}=\sum_{l=0}^{k}{\textstyle\binom{k}{l}}H(\lambda)^{l}(E_{\lambda}^{(K)})^{k-l}.$ (71)

### v3

@@ Appendix C Details of the computer algebra for general spin-1/2 systems
We discuss the details of the computer algebra algorithm and analyze its computational time. We focus entirely on spin-$1/2$ systems for concreteness, while a similar analysis should be possible for spin systems with any spin quantum number and fermionic systems.
@@
We propose a computer algebra framework of spin-$1/2$ systems. The idea of using algebraic computations in the quantum mechanics of spin-$1/2$ systems is not new; it has appeared repeatedly, but rather sporadically, in the literature [76, 77, 78, 72, 33, 79]. Our framework inherits the basic concepts from these previous frameworks, but is more efficient and better suited for our purpose.
@@
is an operator composed of three terms. We use $\sigma,\tau,\dots$ to denote the tensor products of Pauli operators, including $I$ as a special case. We define $z_{A,\sigma}$ as the scalar coefficient before $\sigma$ in $A$. For example, $z_{A,\sigma}=1.5$ for $\sigma=X_{1}Z_{3}$ in Eq. (60). We define $|A|$ as the number of terms in $A$, e.g., $|A|=3$ in the example in Eq. (60). Below, we assume that all terms are $k$-local, with $k$ small and independent of the system size. Based on this assumption, we will disregard the number of Pauli operators per term and only count the number of terms to evaluate the computational time.
Our computer algebra framework stores an operator using a data structure called a hash table [89]. A hash table is a set of correspondences from a “key” to a “value,” which, in our case, is from a tensor product $\sigma$ to the scalar coefficient $z_{A,\sigma}$. A tensor product of Pauli operators is represented by a list of Pauli operator kinds and indices, sorted in ascending order of the spin index. For example, $X_{1}Z_{3}$ is represented by a list $[(X,1),(Z,3)]$, and the entire operator $A$ in Eq. (60) is stored as
$\bigl\{$ $[(X,1),(Z,3)]\mapsto 1.5,\quad[(X,2),(Y,3),(Y,4)]\mapsto-0.3,$
@@
A hash table allows a quick lookup of its elements [89]. Given an arbitrary $\sigma$ and a hash-table representation of an arbitrary operator $A$, it only takes $O(1)$ time to determine whether $\sigma$ is contained in $A$ and, if so, to obtain the scalar coefficient $z_{A,\sigma}$. This search time is independent of $|A|$ for ideal hash tables.
@@
The multiplication of two operators, $A$ and $B$, is completed in $O(|A||B|)$ time by the following procedure. We prepare an empty hash table for storing the result. We pick up one term $\sigma\mapsto z_{A,\sigma}$ from $A$ and one term $\tau\mapsto z_{B,\tau}$ from $B$, multiply the two tensor products $\sigma$ and $\tau$ according to the rules such as $X_{i}Y_{i}=iZ_{i}$, $Y_{i}Z_{i}=iX_{i},$ $Z_{i}X_{i}=iY_{i}$ and $X_{i}X_{i}=Y_{i}Y_{i}=Z_{i}Z_{i}=I_{i}$, multiply their scalar coefficients, and insert $\sigma\tau\mapsto z_{A,\sigma}z_{B,\tau}$ into the hash table for storing the result. This procedure is done in $O(1)$ time. We repeat this procedure for all pairs of terms. Since there are $|A||B|$ pairs, the total computation time is $O(|A||B|)$. This computation time is also expressed as $O(|AB|)$ because the resulting product $AB$ usually contains $|AB|=O(|A||B|)$ terms.
@@
We can efficiently compute traces and commutators by special algorithms that take less time than naive approaches. These algorithms will be crucial for efficiently calculating weighted actions.
The trace $\mathop{\mathrm{tr}}(AB)$ for two operators $A$ and $B$ can be computed in $O(\min\{|A|,|B|\})$ time, which is faster than computing the product $AB$. To achieve this performance, we use the fact that $\mathop{\mathrm{tr}}(\sigma)=0$ for any tensor product of Pauli operators $\sigma\neq I$, except that $\mathop{\mathrm{tr}}(I)=2^{N}$. Moreover, the product $\sigma\tau$ becomes the identity operator if and only if $\sigma=\tau$. Therefore, the trace $\mathop{\mathrm{tr}}(AB)$ is given by $\mathop{\mathrm{tr}}(AB)=2^{N}\sum_{\sigma}z_{A,\sigma}z_{B,\sigma}$, where the sum is over all $\sigma$’s that appear in both $A$ and $B$. If $|A|\leq|B|$, we can calculate $2^{-N}\mathop{\mathrm{tr}}(AB)$ by picking up each term $\sigma\mapsto z_{A,\sigma}$ in $A$, looking up whether $\sigma$ appears in $B$, and if so, adding $z_{A,\sigma}z_{B,\sigma}$ to the result. The single lookup process is done in $O(1)$ time regardless of $|B|$ using a hash table, and thus, the total computational time is $O(|A|)$. If $|A|\geq|B|$, we can swap the roles of $A$ and $B$ to complete the calculation in $O(|B|)$ time. Therefore, we can compute $\mathop{\mathrm{tr}}(AB)$ in $O(\min\{|A|,|B|\})$ time.
@@
The commutator $[B,A]$ for two operators $B$ and $A$ also has a fast algorithm. The commutator of two terms is zero if they have no spins in common, e.g., $X_{1}Y_{2}Z_{3},X_{4}Y_{5}]=0$. Therefore, we do not need to iterate through all pairs of terms of $B$ and $A$, but we can skip most of them. This shortcut is achieved by the following procedure. First, we assign integers $1,2,\dots,|A|$ to the terms of $A$ in an arbitrary order, and we prepare the set of indices $\Theta_{i}(A)=\{l\mid\text{$l$th term of $A$ acts on spin $i$}\}$ for $i=1,\dots,N$. This is done by iterating through the terms of $A$ once, taking $O(|A|)$ time. Next, we prepare an empty hash table to store the result of $[B,A]$. We then iterate through each term $\sigma\mapsto z_{B,\sigma}$ of $B$. If, for example, $\sigma$ acts on spins $i=3,4$, we compute the commutator between $\sigma$ and the terms of $A$ whose indices are in $\Theta_{3}(A)\cup\Theta_{4}(A)$, and we add them to the hash table for storing the result.
This procedure for computing $[B,A]$ runs in $O(N^{-1}|A||B|)$ time under the following assumptions in addition to the $k$-local assumption for all operators. We assume that $A$ consists of $O(N)$ or more terms, and the terms are spread evenly over an $O(N)$ number of spins so that each spin is acted by at most $O(N^{-1}|A|)$ terms. Thus, the size of the index set $|\Theta_{i}(A)|$ is at most $O(N^{-1}|A|)$. For every term in $B$, we compute the commutator with the union of at most $k$ sets out of $\Theta_{1}(A),\dots,\Theta_{N}(A)$ because $B$ is a $k$-local operator. The union contains at most $O(N^{-1}|A|)$ terms, where we neglect $k$ because $k$ is assumed to be small. Thus, the iteration through $B$ takes $O(N^{-1}|A||B|)$ time in total. This analysis also shows that the number of terms in the resulting commutator $|[B,A]|$ is at most $O(N^{-1}|A||B|)$.
By extending this algorithm, we can also efficiently compute the commutators $B,A_{\mu}]$ between an operator $B$ and a series of operators $A_{1},A_{2},\dots$ simultaneously. To do so, we prepare the set $\Theta_{i}(\{A_{\mu}\})=\{(\mu,l)\mid\text{$l$th term of $A_{\mu}$ acts on spin $i$}\}$, and we compute the commutators $B,A_{1}],[B,A_{2}],\ldots$ simultaneously by iterating through the terms of $B$ once, similarly to the above procedure. The computation time is estimated by replacing $|A|$ with $\sum_{\mu}|A_{\mu}|$ in the previous paragraph, assuming that $A_{1},A_{2},\dots$ have $O(N)$ or more terms in total, and the terms spread over $O(N)$ number of spins without concentrating on a few spins. Thus, the total computation takes $O(N^{-1}|B|\sum_{\mu}|A_{\mu}|)$ time, and the resulting number of terms $\sum_{\mu}|[B,A_{\mu}]|$ is at most $O(N^{-1}|B|\sum_{\mu}|A_{\mu}|)$.
@@
We detail the entire algorithm for computing the optimal driving coefficients $\bm{\alpha}^{(\mathcal{P})}(\lambda)$ in the general framework in Sec. II and evaluate its computation time. In computational time analysis, we only count the power of $N$ and neglect prefactors that depend mildly on $K$, since the values of $K$ considered in this paper is small. We use notation from Appendices C.1 and C.2: the symbol $|A|$ denotes the number of terms in $A$, and we use $\mathop{\mathrm{tr}}(A*B)$ to emphasize that we evaluate the trace by preparing $A$ and $B$ without computing the entire product $AB$.
@@
$\partial{}_{\lambda}[H(\lambda)^{k}]=\sum_{l=0}^{k-1}H(\lambda)^{l}[\partial_{\lambda}H(\lambda)]H(\lambda)^{k-1-l}.$ (62)
@@
(3) Compute the fictitious Hamiltonian $\mathcal{P}_{\lambda}(H(\lambda))$ and its derivative $\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))$,
$\mathcal{P}_{\lambda}(H(\lambda))$ $=\sum_{k=0}^{K}p_{k}(\lambda)H(\lambda)^{k},$ (63a)
$\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))$ $=\sum_{k=1}^{K}p_{k}(\lambda)\partial{}_{\lambda}[H(\lambda)^{k}$], (63b)
@@
(4) Compute the commutators $\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]$ simultaneously for all $\mu$. Since $|\mathcal{P}_{\lambda}(H(\lambda))|=O(N^{K})$ and $\sum_{\mu}|A_{\mu}|=O(N)$, the computation time is
$O\biggl(N^{-1}|\mathcal{P}_{\lambda}(H(\lambda))|\sum_{\mu}|A_{\mu}|\biggr)=O(N^{-1}N^{K}N)=O(N^{K}),$ (64)
and the resulting commutators contain $O(N^{K})$ terms in total.
(5) For each $\mu$, compute the commutators $\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}],A_{\nu}]$ for $\nu=1,\dots,\mu$ simultaneously. Using $\sum_{\mu}|[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]|=O(N^{K})$ and $\sum_{\nu=1}^{\mu}|A_{\nu}|=O(N)$ for large enough $\mu$, the total computation time is
$\sum_{\mu}O\biggl(N^{-1}|[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]|\sum_{\nu:\,\nu\leq\mu}|A_{\nu}|\biggr)\qquad$
$=O(N^{-1}N^{K}N)=O(N^{K}),$ (65)
and the total number of resulting terms is $O(N^{K})$.
(6) Compute $r_{\mu}^{(\mathcal{P})}(\lambda)$ for all $\mu$ by
$r_{\mu}^{(\mathcal{P})}(\lambda)=i\hbar\,\mathop{\mathrm{tr}}\left\{\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))*[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]\right\}.$ (66)
Since $|[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]|$ is less than $|\partial^{\prime}_{\lambda}\mathcal{P}(H(\lambda))|$, the calculation of $r_{\mu}^{(\mathcal{P})}(\lambda)$ for a single $\mu$ is done in $O\bigl(|[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]|\bigr)$ time. Thus, the calculation for all $\mu$ takes time $\sum_{\mu}O\bigl(|[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}]|\bigr)=O(N^{K})$.
(7) Compute $Q_{\mu\nu}^{(K)}(\lambda)$ for all $(\mu,\nu)$ with $\mu\geq\nu$ by
$Q_{\mu\nu}^{(\mathcal{P})}(\lambda)=\mathop{\mathrm{tr}}\left\{\mathcal{P}_{\lambda}(H(\lambda))*[[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}],A_{\nu}]\right\},$ (67)
where we use $\mathop{\mathrm{tr}}([A,B]C)=\mathop{\mathrm{tr}}(A[C,B])$ to deduce Eq. (67) from Eq. (17a). Since $|[[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}],A_{\nu}]|$ is less than $|\mathcal{P}_{\lambda}(H(\lambda))|$, the calculation for a single $(\mu,\nu)$ takes $O(|[[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}],A_{\nu}]|)$ time. Thus, the calculation for all $(\mu,\nu)$ with $\mu\geq\nu$ takes the time $\sum_{\mu,\nu:\,\mu\geq\nu}O\bigl(|[[\mathcal{P}_{\lambda}(H(\lambda)),A_{\mu}],A_{\nu}]|\bigr)=O(N^{K})$. Note that $Q_{\mu\nu}^{(K)}(\lambda)$ for $\mu<\nu$ is known from the symmetry $Q_{\mu\nu}^{(K)}(\lambda)=Q_{\nu\mu}^{(K)}(\lambda)$.
(8) Solve the $M$-variate linear equation in Eq. (18) to determine $\bm{\alpha}^{(\mathcal{P})}(\lambda)$. Output $\bm{\alpha}^{(\mathcal{P})}(\lambda)$.
In total, the time for the algebraic computation (steps (1)–(7)) scales as $O(N^{K})$. Step (8) takes $O(M^{3})$ or $O(RM^{2})$ time depending on the numerical algorithm, as discussed in the main text (see Sec. II.4).
Algorithm 2.—This algorithm assumes that the Hamiltonian has the form of Eq. (19). We use symbols and definitions in Appendix A.4.
@@
(1) Compute the operators $\tilde{F}_{g}$ for all $g$ by multiplication and addition of $F_{\gamma}$’s.
(2) For each $g$, compute the commutators $\tilde{F}_{g},A_{\mu}]$ simultaneously for all $\mu$.
(3) For each $g$ and $\mu$, compute the commutators $\tilde{F}_{g},A_{\mu}],A_{\nu}]$ for $\nu=1,\dots,\mu$ simultaneously.
(4) For all $(g,g^{\prime})$, compute
@@
(5) Compute the scalar coefficient $\tilde{f}_{g}(\lambda)$ for all $g$ by multiplying $f_{\gamma}(\lambda)$’s, and compute $\partial_{\lambda}\tilde{f}_{g}(\lambda)$ for all $g$ using the Leibniz rule.
(6) Compute $r_{\mu}^{(\mathcal{P})}(\lambda)$ for all $\mu$ and $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ for all $(\mu,\nu)$ with $\mu\geq\nu$ by Eq. (46). Note that $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ for $\mu<\nu$ is known from the symmetry $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)=Q_{\nu\mu}^{(\mathcal{P})}(\lambda)$.
@@
The computation complexity for the first stage, which deals with algebraic computation, scales as $O(N^{K})$. This is understood by analyzing the computation time for each step similarly to the corresponding step in Algorithm 1. Concretely, step (1) in Algorithm 2 corresponds to step (1) in Algorithm 1, (2) corresponds to (4), (3) corresponds to (5), and (4) corresponds to (6) and (7).
The computation cost of the second stage, which involves only scalar computation, is dominated by step (7), requiring $O(M^{3})$ or $O(RM^{2})$ time depending on the numerical algorithm. Step (5) takes $O(M^{0})$ time, and step (6) takes $O(M^{2})$ time, which are faster than step (7).
@@
This subsection presents entire algorithms specialized for ground-state evolution in Sec. III. These algorithms combine the general algorithms in Appendix C.3 with the minimization of $\Omega_{\lambda}^{(K)}(E)$ discussed in Appendices B.2 and B.3. The general algorithms are slightly simplified by taking advantage of the simple form of the polynomial, $\mathcal{P}^{\mathrm{GS},K}(H(\lambda))=[H(\lambda)-E_{\lambda}^{(K)}]^{K}$. As in Appendix C.3, we present two algorithms, algorithm 3 for a single value of $\lambda$ and algorithm 4 for an entire range of $\lambda$. Our C++ code [73] implements algorithm 3 and algorithm 4.
Algorithm 3.—For a fixed $\lambda$, this algorithm takes the algebraic representations of $H(\lambda)$, $\partial_{\lambda}H(\lambda)$, and $A_{\mu}$ for $\mu=1,\dots,M$ as input. The algorithm proceeds as follows:
@@
(3) Minimize the function $\Omega_{\lambda}^{(K)}(E)$ to determine $E_{\lambda}^{(K)}$. The details of the minimization are discussed in Appendix B.2.
(4) Compute $H(\lambda)-E_{\lambda}^{(K)}]^{k}$ for $k=2,\dots,K$ by
$H(\lambda)-E_{\lambda}^{(K)}]^{k}=\sum_{l=0}^{k}{\textstyle\binom{k}{l}}H(\lambda)^{l}(-E_{\lambda}^{(K)})^{k-l}.$ (70)